### v2

# Karabo: A versatile SKA Observation Simulation Framework

Abstract
Karabo is a versatile Python-based software framework simplifying research with radio astronomy data. It bundles existing software packages into a coherent whole to improve the ease of use of its components. Karabo includes useful abstractions, like strategies to scale and parallelize typical workloads or science-specific Python modules. The framework includes functionality to access datasets and mock observations to study the Square Kilometer Array (SKA) instruments and their expected accuracy. SKA will address problems in a wide range of fields of astronomy. We demonstrate the application of Karabo to some of the SKA science cases from HI intensity mapping, mock radio surveys, radio source detection, the epoch of re-ionisation and heliophysics. We discuss the capabilities and challenges of simulating large radio datasets in the context of SKA.
keywords:
[…]
[third]organization=Institute for Particle Physics and Astrophysics, ETH Zurich,addressline=Wolfgang-Pauli-Strasse 27, city=Zurich, postcode=8093, country=Switzerland
[fourth]organization=Laboratoire d’Astrophysique, Ecole Polytechnique Federale de Lausanne EPFL, adressline=Observatoire de Sauverny, city=Versoix, postcode=1290, country=Switzerland
1 Introduction
Radio observations are crucial for the understanding of the universe. They provide a unique window to the universe, and the upcoming world’s largest radio interferometer, the Square Kilometer Array (SKA), will provide unprecedented radio data quality to enable data-driven discoveries. However, in any data-driven discovery, the simulations of radio observations play a crucial role in providing a comprehensive picture in most radio science cases. They quantify the instrumental artefacts and relevance of the desired radio signal and debug the data pipelines.
SKA is currently under construction (Dewdney et al., 2009). It consists of two instruments, SKA-Low and SKA-Mid, which cover the 80-350 MHz and 350 MHz-30 GHz frequency range, respectively. SKA will enable a vast range of science cases (Braun et al., 2015), with unique sets of scientific and data requirements, hence a challenge. Most SKA measurements will be distributed as 3D data cubes (a 2D sky measured in multiple frequencies). In rare cases, visibilities and calibration solutions will be made available as well. This presents a challenge in debugging the data signal chains. The large quantity of data presents challenges not only to data pipelines but also to data handling and analysis at the user’s end. The users must be able to use their own pieces of code for analysis. The analysis tools must be user-friendly enough to appeal to a diverse range of users with different expertise. Mock radio data generation is an integral and common part of analysis tools.
We have developed Karabo111https://github.com/i4Ds/Karabo-Pipeline, a user-friendly and versatile radio data simulator. The primary motivator of Karabo is SKA, allowing mock simulations from stations and antennas for SKA-low and SKA-mid. We utilise OSKAR222https://github.com/OxfordSKA/OSKAR and RASCIL333https://gitlab.com/ska-telescope/external/rascil packages as simulation engines for SKA-low and SKA-mid, respectively. In Karabo, the distribution of existing radio astronomy software packages is made interoperable through common Python interfaces. Individual functions from existing packages can be combined in a number of ways to build radio-astronomy data processing pipelines for a wide variety of use cases. Karabo provides easy access to several real-world and synthetic datasets. In the following chapters, we outline the Karabo architecture and design choices and present science cases using Karabo.
The scale of the SKA data products exceeds that of all previous instruments (Quinn et al., 2015). To handle the enormous datasets, novel methods have to be developed, and existing algorithms must be adapted. Some of these novel methods might be adapted from other domains dealing with large datasets. For example, machine learning methods can be used to localize sources without imaging (Taran et al., 2023) or for fast image deconvolution (Drozdova et al., 2024). One of our goals is to make it easier for non-domain experts to work with radio astronomy data, algorithms, software, and challenges.
Karabo makes it easy to install and use all relevant software packages. Users can build intricate pipelines in just a few lines of Python code. In addition to SKA, Karabo also supports various existing telescopes, including SKA-precursors and pathfinders.
1.1 Related Work
CASA (Bean et al., 2022) served as inspiration for Karabo, as it follows similar ideals. It bundles many interoperable algorithms in a user-friendly package. Compared to Karabo, CASA is more opinionated in its data models, and more geared towards domain experts. Having a single, standardized data model improves usability and consistency and simplifies the implementation, but it also comes with its own challenges. For example, the reliance on the Measurement Set visibility format limits scalability (Perkins et al., 2021).
We wrap existing software packages with common Python interfaces for interoperability. The interfaces provide a common abstraction for different implementations, e.g., imagers. Users are free to combine the packages into complex data processing pipelines, as shown in Fig. 1. The following chapters describe the fundamental components of Karabo.
1.2 SkyModel
The SkyModel entity defines the sources in the sky, which can be observed in later stages. We define a sky as a combination of point sources and extended sources. The minimum information needed per source is at least Stokes I/Q/U/V, but Karabo can manage additional information, such as Full-Width at Half Maximum (FWHM) or spectral index. We use Jansky as a common unit for the Stokes parameters or input flux density. A SkyModel can be created from code, loaded from well-known catalogues like GLEAM Hurley-Walker et al. (2017) and MALS Gupta et al. (2017), or read from .h5 or catalog files. An example of a small catalog file is shown in Fig. 2. Users can further visualize a SkyModelor filter sources by frequency, flux and position. The supported functions should feel familiar to users coming from OSKAR or RASCIL.
Not every interferometer simulation supports both extended sources or a large number of point sources. We convert between those representations as required, e.g. extended sources may be approximated by a large number of point sources. Similarly, Karabo may bucket continuous spectra into fixed frequency bins when required.
| (A) Skymodel Image | (B) Dirty image | (C) PB corrected Dirty image |
(C) Mosaic for 3 pointings
Fig. 3 (A) shows the flux density of the input sky restricted to the same area as the simulated pointing (Fig. 3 (B)) and plotted with healpix 444http://healpix.sourceforge.net. Using Karabo’s HDF5 file interface, we import the map and build a sky model with the centre at R.A. , and DEC and a diameter of . The simulation was done for the MeerKAT configuration between 700 MHz and 800 MHz with a bandwidth of 5 MHz, and with a primary beam’s full width at half maxima (FWHM) of 1.8o. Fig. 3 (B) shows the dirty image, followed by the primary beam corrected image (C) and a mosaic of 3 pointings (D). Common emission features between the dirty image and the input sky model can be observed (A). The resulting mosaic has a size of . The dirty image and mosaic shown are added up along the redshift axis. However, the simulation also provides three-dimensional images in the chosen redshift resolution.
| (A) Input MIGHTEE skymodel | (B) Mock image from MeerKAT |
| (C) Input GLEAM skymodel | (D) Mock image from MWA |
1.3 Mock Continuum Survey
Radio continuum observations play an essential role in various science cases, for example relating to galaxy studies or cosmology. A radio continuum is a continuous emission in frequency space that arises from multiple types of objects, like active galactic nuclei (AGN), radio lobes, star-forming regions, diffuse thermal and gyrosynchroton emission, etc. This section demonstrates simple mock data simulation for a limited MeerKAT International GHz Tiered Extragalactic Exploration (Jarvis et al., 2016, MIGHTEE) and GaLactic and Extragalactic All-sky MWA survey (Wayth et al., 2015, GLEAM) point source catalogs. These were observed by MeerKAT and MWA, the SKA precursors.
1.3.1 Mock MIGHTEE catalog image
The MIGHTEE survey by the MeerKAT telescope is one of the modern surveys covering a 20 field-of-view (FoV) over the COSMOS, E-CDFS, ELASIS-S1 and XMM-Newton Large Scale Structure field deep fields at SKA-mid frequencies. Using the point source catalogue from MIGHTEE over a COSMOS FoV, we simulate observations from the MeerKAT telescope configuration, which has 64 antennas. The pointing centre was at R.A. = , DEC = . Within the FoV, the catalogue has sources with a distribution in the flux densities with their respective spectral indices. We demonstrate a snapshot simulation, which consists of an 8-second time integration and 0.2 MHz channel bandwidth integration. Measurement sets were created for the total observation was done between 856 MHz and 1675.2 MHz with 4096 frequency channels. The clean images from the measurement set of the observation were created, followed by imaging using WSCLEAN deconvolution. The point sources from the catalogue and the simulated deconvolved image are shown in Fig. 4 (A) & (B).
1.3.2 Mock-GLEAM catalog image
In SKA-low frequencies, the MWA’s GLEAM is the most comprehensive wide-field survey covering 24,831 . This large FoV was achieved by drift-scanning using MWA’s fine 40 kHz frequency channel width and 0.5 sec time resolution. GLEAM contains more than 0.3 million components. We use a subset of these sources in our simulation to demonstrate the mock data creation at SKA-low frequencies using MWA phase-I configuration. We use R.A. and DEC of the pointing at and , respectively. We limit the sky model FoV to around the pointing centre, with 23600 radio sources. We performed a snapshot observation for demonstration, covering an observation duration of 1 minute with 0.5-second time integration and 40 kHz channel bandwidth.
Fig. 4 (C) shows the input GLEAM sky to Karabo, while panel (D) shows the simulated snapshot images from the MWA telescope. As expected, we notice the brightest regions in the simulated images corresponding to the input model (panels A & C).
1.4 Source Detection
Many problems, ranging from galactic and extragalactic studies to cosmology, desire the detection of faint radio point sources. All science cases that build on source statistics and source characterisation are based on source detection. The objective of source detection is to detect radio sources with a given significance level and for each source determine additional parameters like size, flux densities, and spectral index. Karabo performs source detection with PyBDSF (Mohan and Rafferty, 2015). This is useful by itself, but Karabo can also be used to evaluate the accuracy of source detection algorithms, thanks to built-in evaluation metrics like precision and recall. Karabo can compare the found sources with known sources from a catalog. A demonstration of the source detection from Karabo is shown in Fig. 5. Here, we simulated 8 point sources to be detected. Using the ASKAP telescope configuration (Fig. LABEL:fig:gleam-askap (B)), we simulate visibilities at starting frequency 100 MHz, channel bandwidth 1 MHz, 16 frequency channels and 24 time channels. We used the RASCIL imager for the dirty image and WSCLEAN for a cleaned version. The cleaned image was passed to PYBDSF, and source detection was performed (Fig. 5).
1.5 Epoch of Reionisation
After the big bang, the recombination of hydrogen atoms resulted in an 21cm hydrogen line emission. Later, as the first stars formed in our universe, the resulting heating changed the absorption profile of this emission. The resulting subtle changes with redshift or observing frequency is called the Epoch of Reionization (EoR) signal (EoR Zaroubi, 2013).
EoR measurements are challenging due to the weak and extended nature of the signal, the presence of strong and diverse foregrounds, ranging from nearby strong point sources, extended galactic synchrotron emission, weak extra-galactic point source populations, and others. Karabo can simulate EoR signals with various foreground signals making it a great tool for EoR simulation studies. Major foreground signals that can be used are GLEAM catalogue point sources, extended Haslam synchrotron maps, and others. The tools21cm package (Giri et al., 2020) is integrated into Karabo to simulate the EoR signal itself. The combined data offers an exciting glimpse into the future EoR studies that SKA makes possible.
Fig. 6 (A) and (B) shows EoR signal and signal noise simulated from tools21cm modules integrated into the Karabo package. As a part of SKA Data Challenge 3a, the SKACH team used Karabo simulations to study EoR detection and the role of the foreground in the images seen through the SKA-mid configuration (Bonaldi et al., 2025; Bianco et al., 2024). The primary goal of the data challenge was to remove the foreground and recover the EoR signal. Panels (C) and (D) show the galactic foreground and bright point sources, respectively. The bright point sources produce strong sidelobe artefacts which must be removed to recover the faint EoR signal. The point source removal is more effective in visibility space than in the image plane (Sharma et al., 2022). We integrated such a visibility point source removal algorithm into Karabo.
1.6 Heliophysics
The Sun is the nearest and brightest radio source, and most important to human space-based activity. The energetic particle beams originating in the corona are the driver for space weather and need to be understood in terms of their time and frequency variability (e.g. Harra et al., 2021). Radio observations of the sun increase our understanding of the corona, the outermost layer of the sun, and the interplanetary medium. The high sensitivity of the SKA would provide solar images at an unprecedented time and frequency resolution (Nindos et al., 2019; Oberoi et al., 2023). Using Karabo, we can estimate the expected levels of detail of SKA images for different activity levels of the Sun.
Fig. 7 shows an example of the simulation of a solar disk for a snapshot (8-sec time integration). The sky model was obtained by the FORWARD model (Gibson et al., 2016), which simulates the radio emission from thermal bremsstrahlung based on realistic magnetohydrodynamic models from magnetic field maps. We note that the SKA-mid and SKA-low simulated images have good dynamic range. The SKA-low image is more diffuse than SKA-mid image at high frequencies. Many types of such solar simulation studies can be done using Karabo, like simulation of spatio-temporal features of active regions, level of self-noise in the image, and many others. Also, the effect of sidelobes produced by the Sun in non-solar observations can be simulated.
| (A) FORWARD solar skymodel | (B) SKA-mid image | SKA-low image |
2 Discussion & Summary
SKA data simulations give first-hand insights into the upcoming instrument capabilities and challenges. The development and operations of SKA instrumentation are challenges in themselves (Norris, 2011; An, 2019). In addition, radio data analysis of a wide range of science topics is non-trivial and requires simulation support. Simulating a wide range of science cases using common software presents unique capabilities and challenges. The primary challenge is to incorporate multiple interactions of radio waves from celestial radio sources with the telescope configurations, radio frequency environment, ionosphere and instrumental systematics, followed by producing realistic 3-D data cubes. Instrumental systematics is a long list of effects known in radio telescope operations that result in complex changes in the UV plane.
Karabo is an open-source software designed to simulate radio datasets for a wide range of science cases. It tackles all the mentioned challenges systematically. In this study, we demonstrated some of the simulations of a subset of science topics, i.e. mock continuum survey, point source detection, EoR signal and simulation of radio Sun useful for cosmology, transients, EoR and heliophysics science. Some of Karabo’s unique capabilities are as follows:
1.
Karabo’s framework allows both SKA-low and SKA-mid science, i.e. useful for science cases requiring both telescopes in frequency and time.
2.
Many software tooling focuses on a single science case, which Karabo combines to support a broader scope.
3.
Karabo includes relevant metadata such as telescope configurations, and also data sets like sky surveys.
4.
End-to-end implementation of radio data analysis starting from sky models, processing of visibilities, to deconvolution, followed by image data analysis.
5.
Karabo scales from personal computers to HPC clusters and automatically makes use of the available hardware.
As SKA-low and SKA-mid sites cover separate frequency bands, the science cases featuring the frequency band edges can be studied in a friendly way using Karabo. In addition, the location of the low and mid telescopes are far apart and lie in GMT+8 and GMT+2 time zones, respectively. The time difference of 6 hours allows for common time windows for simultaneous observations. Karabo can be used to design such observations and study realistic problems in imaging the desired field of view.
Different fields of radio astronomy follow their software developments. The tools developed are necessary for the overall data analysis and have specific features dedicated to the radio astronomy field. The challenge originates from focused developments with limited interfaces. The interfaces in Karabo support multiple implementations for different science cases. In addition, Karabo also interfaces with different data formats. Depending on the need of the user, software integration can also occur at the data analysis end, i.e., analysing the images and data cubes as demonstrated by source detection in the section.
Karabo supports multiple telescope configurations, including SKA precursor and pathfinder telescopes. This can be used to mimic already working telescopes and preparing workflows involving real data from precursors and pathfinder telescopes. Karabo is an excellent tool for building mock data for source detection, such as studies and machine learning applications.
Karabo is a continuing development aimed at systematically simulating SKA datasets in sync with the development of different array assemblies. The flexible architecture of Karabo also allows implementing the instrumental systematics from actual measurements from SKA’s array assemblies. The inclusion of existing and custom telescope configurations, along with the inclusion of realistic effects, makes Karabo not only interesting for SKA simulations but a powerful, general tool to work with radio observations.
Acknowledgements
This work was supported in part by the AstroSignals Sinergia grant No. CRII5_193826 from the Swiss National Science Foundation. This work was supported by SERI as part of the SKACH consortium. We also thank Christoph Vögele in the initial stages of the project. This work was supported by a grant from the Swiss National Supercomputing Centre (CSCS) under project ID sk05. Some of the results in this paper have been derived using the healpy and HEALPix packages.